I am trying to help my son do 5 step method equations.
The question is
A large soda costs $1 more than a hamburger. If 4 hamburges and 2 sodas cost $14, what is the cost of each? Show all steps.
Can someone help me so I know how to help him with the rest of his problems. Thanks, Kami

Answer by aaronwiz(69)

Hey Kami,
My name is Aaron, and I am in 10th grade and am in honors trig and pre calc. I feel for you, and will do my best to help. What your son needs to do is set up a system of equations.
Let s=large soda
and h=hamburger
first "A large soda costs $1 more than a hamburger" so...
h+1=s
next "4 hamburges and 2 sodas cost $14" so...
4h+2s=14
put it together, and you have a system of equations...
h+1=s
4h+2s=14
I'm sure by now either you or your son can solve this, but just to be safe il finish it.
Their are two ways to go about this either substitution or combinations/elimination method
lets use substitution since we already have s=h+1
substitute this into s for the second equation
4h+2(h+1)=14
4h+2h+2=14
6h=12
h=2
plug this back into one of the original equations, ill use h+1=s
(2)+1=s
so s=3
thus a hamburger is $2 and a soda is $3
Hope I helped, good luck to you and your son. Positive feedback is always appreciated.